# GaLore: Memory-Efficient LLM Training by Gradient Low-Rank Projection

### v1

@@ 4.1 Composition of Low-Rank Subspaces
Figure 2: Learning through low-rank subspaces $\Delta W_{T_{1}}$ and $\Delta W_{T_{2}}$ using GaLore. For $t_{1}\in[0,T_{1}-1]$, $W$ are updated by projected gradients $\tilde{G}_{t_{1}}$ in a subspace determined by fixed $P_{t_{1}}$ and $Q_{t_{1}}$. After $T_{1}$ steps, the subspace is changed by re-computing $P_{t_{2}}$ and $Q_{t_{2}}$ for $t_{2}\in[T_{1},T_{2}-1]$, and the process repeats until convergence.
@@
Following the above procedure, the switching frequency $T$ becomes a hyperparameter. The ablation study (Fig. 5) shows a sweet spot exists. A very frequent subspace change increases the overhead (since new $P_{t}$ and $Q_{t}$ need to be computed) and breaks the condition of constant projection in Theorem 3.6. In practice, it may also impact the fidelity of the optimizer states, which accumulate over multiple training steps. On the other hand, a less frequent change may make the algorithm stuck into a region that is no longer important to optimize (convergence proof in Theorem 3.6 only means good progress in the designated subspace, but does not mean good overall performance). While optimal $T$ depends on the total training iterations and task complexity, we find that a value between $T=50$ to $T=1000$ makes no significant difference. Thus, the total computational overhead induced by SVD is negligible ($<10\%$) compared to other memory-efficient training techniques such as memory offloading (Rajbhandari et al., 2020).
@@ 4.2 Memory-Efficient Optimization
Reducing memory footprint of gradient statistics. GaLore significantly reduces the memory cost of optimizer that heavily rely on component-wise gradient statistics, such as Adam (Kingma & Ba, 2014). When $\rho_{t}\equiv\mathrm{Adam}$, by projecting $G_{t}$ into its low-rank form $R_{t}$, Adam’s gradient regularizer $\rho_{t}(R_{t})$ only needs to track low-rank gradient statistics. where $M_{t}$ and $V_{t}$ are the first-order and second-order momentum, respectively. GaLore computes the low-rank normalized gradient $N_{t}$ as follows:
@@ 4.3 Combining with Existing Techniques
GaLore is compatible with existing memory-efficient optimization techniques. In our work, we mainly consider applying GaLore with 8-bit optimizers (Dettmers et al., 2021) and per-layer weight updates (Lv et al., 2023).
@@
Dettmers et al. (2022) proposed 8-bit Adam optimizer that maintains 32-bit optimizer performance at a fraction of the original memory footprint. We apply GaLore directly to the existing implementation of 8-bit Adam.
@@
In practice, the optimizer typically performs a single weight update for all layers after backpropagation. This is done by storing the entire weight gradients in memory. To further reduce the memory footprint during training, we adopt per-layer weight updates to GaLore, which performs the weight updates during backpropagation (Lv et al., 2023).
@@ 5 Experiments
Table 2: Comparison with low-rank algorithms on pre-training various sizes of LLaMA models on C4 dataset. Validation perplexity is reported, along with a memory estimate of the total of parameters and optimizer states based on BF16 format. The actual memory footprint of GaLore is reported in Fig. 1 and Fig. 4.
@@
We evaluate GaLore on both pre-training and fine-tuning of LLMs. All experiments are conducted on NVIDIA A100 GPUs [^2] The implementation of GaLore is available here.
@@
To evaluate its performance, we apply GaLore to train LLaMA-based large language models on the C4 dataset. C4 dataset is a colossal, cleaned version of Common Crawl’s web crawl corpus, which is mainly intended to pre-train language models and word representations (Raffel et al., 2023). To best simulate the practical pre-training scenario, we train without data repetition over a sufficiently large amount of data, across a range of model sizes up to 7 Billion parameters.
@@
We follow the experiment setup from Lialin et al. (2023), which adopts a LLaMA-based [^3] LLaMA materials in our paper are subject to LLaMA community license. architecture with RMSNorm and SwiGLU activations (Touvron et al., 2023; Zhang & Sennrich, 2019; Shazeer, 2020). For each model size, we use the same set of hyperparameters across methods, except the learning rate. We run all experiments with BF16 format to reduce memory usage, and we tune the learning rate for each method under the same amount of computational budget and report the best performance. The details of our task setups and hyperparameters are provided in the appendix.
@@
5.1 Comparison with low-rank methods
@@
Lialin et al. (2023) is a variant of LoRA designed for pre-training, which periodically merges $BA$ into $W$, and initializes new $BA$ with a reset on optimizer states and learning rate. ReLoRA requires careful tuning of merging frequency, learning rate reset, and optimizer states reset. We evaluate ReLoRA without a full-rank training warmup for a fair comparison.
Mem 40K 80K 120K 150K
8-bit GaLore 18G 17.94 15.39 14.95 14.65
8-bit Adam 26G 18.09 15.47 14.83 14.61
Tokens (B) 5.2 10.5 15.7 19.7
For GaLore, we set subspace frequency $T$ to 200 and scale factor $\alpha$ to 0.25 across all model sizes in Table 2. For each model size, we pick the same rank $r$ for all low-rank methods, and we apply them to all multi-head attention layers and feed-forward layers in the models. We train all models using Adam optimizer with the default hyperparameters (e.g., $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$). We also estimate the memory usage based on BF16 format, including the memory for weight parameters and optimizer states. As shown in Table 2, GaLore outperforms other low-rank methods and achieves comparable performance to full-rank training. We note that for 1B model size, GaLore even outperforms full-rank baseline when $r=1024$ instead of $r=512$. Compared to LoRA and ReLoRA, GaLore requires less memory for storing model parameters and optimizer states. A detailed training setting of each model and our memory estimation for each method are provided in the appendix.
@@ 5.2 GaLore with Memory-Efficient Optimizers
We demonstrate that GaLore can be applied to various learning algorithms, especially memory-efficient optimizers, to further reduce the memory footprint. We apply GaLore to AdamW, 8-bit Adam, and Adafactor optimizers (Loshchilov & Hutter, 2019; Dettmers et al., 2022; Shazeer & Stern, ). We consider Adafactor with first-order statistics to avoid performance degradation.
@@ 5.3 Scaling up to LLaMA 7B Architecture
Scaling ability to 7B models is a key factor for demonstrating if GaLore is effective for practical LLM pre-training scenarios. We evaluate GaLore on an LLaMA 7B architecture with an embedding size of 4096 and total layers of 32. We train the model for 150K steps with 19.7B tokens, using 8-node training in parallel with a total of 64 A100 GPUs. Due to computational constraints, we only compare 8-bit GaLore ($r=1024$) with 8-bit Adam with a single trial without tuning the hyperparameters. As shown in Table 3, after 150K steps, 8-bit GaLore achieves a perplexity of 14.65, which is comparable to 8-bit Adam with a perplexity of 14.61.
@@ 5.5 Measurement of Memory and Throughput
Training 7B models on consumer GPUs with 24G memory. As shown in Fig. 4, 8-bit GaLore requires significantly less memory than BF16 baseline and 8-bit Adam, and only requires 22.0G memory to pre-train LLaMA 7B with a small per-GPU token batch size (up to 500 tokens). This memory footprint is within 24GB VRAM capacity of a single GPU such as NVIDIA RTX 4090. In addition, when activation checkpointing is enabled, per-GPU token batch size can be increased up to 4096. While the batch size is small per GPU, it can be scaled up with data parallelism, which requires much lower bandwidth for inter-GPU communication, compared to model parallelism. Therefore, it is possible that GaLore can be used for elastic training (Lin et al., ) 7B models on consumer GPUs such as RTX 4090s.
@@ 6 Ablation Study
6.1 How many subspaces are needed during pre-training?
We observe that both too frequent and too slow changes of subspaces hurt the convergence, as shown in Fig. 5(left). The reason has been discussed in Sec. 4.1 and is more prevalent for small $r$, since in such case, the subspace switching should happen at the right time to avoid wasting optimization steps in the wrong subspace, while for large $r$ the gradient updates cover more subspaces, providing more cushion.
@@ 6.1 How many subspaces are needed during pre-training?
6.2 How does the rank of subspace affect the convergence?
Within a certain range of rank values, decreasing the rank only slightly affects the convergence rate, causing a slowdown that is close to linear. As shown in Fig. 5(right), training with a rank of 128 using 80K steps achieves a lower loss than training with a rank of 512 using 20K steps. This shows that GaLore can be used to trade-off between memory and computational cost. In a memory-constrained scenario, reducing the rank allows us to stay within the memory budget while training for more steps to preserve the performance.
@@ 7 Conclusion
We identify several open problems for GaLore, which include (1) applying GaLore on training of other types of models such as vision transformers and diffusion models, (2) further improving memory efficiency by employing low-memory projection matrices, through quantization or special parameterization, and (3) exploring the possibility of elastic data distributed training on low-bandwidth consumer-grade hardware.
@@
We hope that our work will inspire future research on memory-efficient LLM training strategies from the perspective of low-rank gradient projection. We believe that GaLore will be a valuable tool for the community to train large language models with consumer-grade hardware and limited resources.
8 Impact Statement
@@
This paper aims to improve the memory efficiency of training large language models (LLMs) in order to reduce the environmental impact of LLM pre-training and fine-tuning. By enabling the training of larger models on hardware with lower memory, our approach helps to minimize energy consumption and carbon footprint associated with training LLMs.
Appendix A Proofs
A.1 Gradient becomes low-rank
See 3.1
Proof.
We have
$G_{t}=A-BW_{t}C=A-B(W_{t-1}+\eta G_{t-1})C=G_{t-1}-\eta BG_{t-1}C$ (18)
Let $B=UD_{B}U^{\top}$ and $C=VD_{C}V^{\top}$ be the eigen decomposition of $B$ and $C$. $D_{B}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{m})$ and $D_{C}=\operatorname{diag}(\nu_{1},\ldots,\nu_{n})$ are their eigenvalues sorted in ascending orders (i.e., $\lambda_{1}\leq\ldots\leq\lambda_{m}$ and $\nu_{1}\leq\ldots\leq\nu_{n}$). Define $H_{t}:=U^{\top}G_{t}V$. It is clear that $\mathrm{rank}(H_{t})=\mathrm{rank}(G_{t})$ and we have:
$H_{t}:=U^{\top}G_{t}V=H_{t-1}-\eta D_{B}H_{t-1}D_{C}$ (19)
Suppose $h_{t,ij}$ is the $ij$ component of $H_{t}$, then from the equation above we have:
$h_{t,ij}=h_{t-1,ij}-\eta\lambda_{i}\nu_{j}h_{t-1,ij}=(1-\eta\lambda_{i}\nu_{j})h_{t-1,ij}=(1-\eta\lambda_{i}\nu_{j})^{t}h_{0,ij}$ (20)
Then for first few rows $i$ and columns $j$ that correspond to large eigenvalues, $h_{t,ij}\rightarrow 0$ quickly and $\mathrm{rank}(H_{t})$ becomes small.
To make it more precise, consider the stable rank:
[…]
Then we have:
$\|H_{t}\|_{F}^{2}=\sum_{i=1}^{m}\sum_{j=1}^{n}(1-\eta\lambda_{i}\nu_{j})^{2t}h^{2}_{0,ij}$ (22)
and
$\|H_{t}\|_{2}^{2}\geq\sum_{j=1}^{n}H^{2}_{t,1j}=\sum_{j=1}^{n}(1-\eta\lambda_{1}\nu_{j})^{2t}h^{2}_{0,1j}$ (23)
With high probability, $h^{2}_{0,1j}\geq\epsilon^{2}_{0}$, since $|h^{2}_{1i}|\leq c_{0}$ is bounded, we have:
[…]
Using Mediant inequality, $\frac{a}{b}\leq\frac{a+c}{b+d}\leq\frac{c}{d}$ for $a,b,c,d>0$, therefore, we know that for $i$-th row ($i\geq 2$), since $\lambda_{i}\geq\lambda_{1}$:
$\frac{\sum_{j=1}^{n}(1-\eta\lambda_{i}\nu_{j})^{2t}}{\sum_{j=1}^{n}(1-\eta\lambda_{1}\nu_{j})^{2t}}\leq\max_{j}\left(\frac{1-\eta\lambda_{i}\nu_{j}}{1-\eta\lambda_{1}\nu_{j}}\right)^{2t}=\left(\frac{1-\eta\lambda_{i}\nu_{1}}{1-\eta\lambda_{1}\nu_{1}}\right)^{2t}\leq 1$ (25)
and the conclusion follows. ∎
A.2 Reversibility
Definition A.1 (Reversiblity (Tian et al., 2020)).
@@ A.2 Reversibility
A network $\mathcal{N}$ that maps input ${\bm{x}}$ to output ${\bm{y}}=\mathcal{N}({\bm{x}})$ is reversible, if there exists $K({\bm{x}};W)$ so that ${\bm{y}}=K({\bm{x}};W){\bm{x}}$, and the backpropagated gradient ${\bm{g}}_{\bm{x}}$ satisfies ${\bm{g}}_{\bm{x}}=K^{\top}({\bm{x}};W){\bm{g}}_{\bm{y}}$, where ${\bm{g}}_{\bm{y}}$ is the backpropagated gradient at the output ${\bm{y}}$. Here $K({\bm{x}};W)$ depends on the input ${\bm{x}}$ and weight $W$ in the network $\mathcal{N}$.
@@
See 3.3

### v2

@@ 4.1 Composition of Low-Rank Subspaces
Figure 2: Learning through low-rank subspaces $\Delta W_{T_{1}}$ and $\Delta W_{T_{2}}$ using GaLore. For $t_{1}\in[0,T_{1}-1]$, $W$ are updated by projected gradients $\tilde{G}_{t_{1}}$ in a subspace determined by fixed $P_{t_{1}}$ and $Q_{t_{1}}$. After $T_{1}$ steps, the subspace is changed by recomputing $P_{t_{2}}$ and $Q_{t_{2}}$ for $t_{2}\in[T_{1},T_{2}-1]$, and the process repeats until convergence.
@@
Following the above procedure, the switching frequency $T$ becomes a hyperparameter. The ablation study (Fig. 5) shows a sweet spot exists. A very frequent subspace change increases the overhead (since new $P_{t}$ and $Q_{t}$ need to be computed) and breaks the condition of constant projection in Theorem 3.8. In practice, it may also impact the fidelity of the optimizer states, which accumulate over multiple training steps. On the other hand, a less frequent change may make the algorithm stuck into a region that is no longer important to optimize (convergence proof in Theorem 3.8 only means good progress in the designated subspace, but does not mean good overall performance). While optimal $T$ depends on the total training iterations and task complexity, we find that a value between $T=50$ to $T=1000$ makes no much difference. Thus, the total computational overhead induced by SVD is negligible ($<10\%$) compared to other memory-efficient training techniques such as memory offloading (Rajbhandari et al., 2020).
@@ 4.2 Memory-Efficient Optimization
Reducing memory footprint of gradient statistics. GaLore significantly reduces the memory cost of optimizer that heavily rely on component-wise gradient statistics, such as Adam (Kingma & Ba, 2015). When $\rho_{t}\equiv\mathrm{Adam}$, by projecting $G_{t}$ into its low-rank form $R_{t}$, Adam’s gradient regularizer $\rho_{t}(R_{t})$ only needs to track low-rank gradient statistics. where $M_{t}$ and $V_{t}$ are the first-order and second-order momentum, respectively. GaLore computes the low-rank normalized gradient $N_{t}$ as follows:
@@ 4.3 Combining with Existing Techniques
GaLore is compatible with existing memory-efficient optimization techniques. In our work, we mainly consider applying GaLore with 8-bit optimizers and per-layer weight updates.
@@
Dettmers et al. (2022) proposed 8-bit Adam optimizer that maintains 32-bit optimizer performance at a fraction of the memory footprint. We apply GaLore directly to the existing implementation of 8-bit Adam.
@@
In practice, the optimizer typically performs a single weight update for all layers after backpropagation. This is done by storing the entire weight gradients in memory. To further reduce the memory footprint during training, we adopt per-layer weight updates to GaLore, which performs the weight updates during backpropagation. This is the same technique proposed in recent works to reduce memory requirement (Lv et al., 2023a, b).
@@ 5 Experiments
Table 2: Comparison with low-rank algorithms on pre-training various sizes of LLaMA models on C4 dataset. Validation perplexity is reported, along with a memory estimate of the total of parameters and optimizer states based on BF16 format. The actual memory footprint of GaLore is reported in Fig. 4.
@@
We evaluate GaLore on both pre-training and fine-tuning of LLMs. All experiments run on NVIDIA A100 GPUs.
Mem 40K 80K 120K 150K
8-bit GaLore 18G 17.94 15.39 14.95 14.65
8-bit Adam 26G 18.09 15.47 14.83 14.61
Tokens (B) 5.2 10.5 15.7 19.7
@@
To evaluate its performance, we apply GaLore to train LLaMA-based large language models on the C4 dataset. C4 dataset is a colossal, cleaned version of Common Crawl’s web crawl corpus, which is mainly intended to pre-train language models and word representations (Raffel et al., 2020). To best simulate the practical pre-training scenario, we train without data repetition over a sufficiently large amount of data, across a range of model sizes up to 7 Billion parameters.
@@
We follow the experiment setup from Lialin et al. (2024), which adopts a LLaMA-based [^3] LLaMA materials in our paper are subject to LLaMA community license. architecture with RMSNorm and SwiGLU activations (Zhang & Sennrich, 2019; Shazeer, 2020; Touvron et al., 2023). For each model size, we use the same set of hyperparameters across methods, except the learning rate. We run all experiments with BF16 format to reduce memory usage, and we tune the learning rate for each method under the same amount of computational budget and report the best performance. The details of our task setups and hyperparameters are provided in the appendix.
@@
5.1 Comparison with Existing Low-Rank Methods
@@
Lialin et al. (2024) proposed ReLoRA, a variant of LoRA designed for pre-training, which periodically merges $BA$ into $W$, and initializes new $BA$ with a reset on optimizer states and learning rate. ReLoRA requires careful tuning of merging frequency, learning rate reset, and optimizer states reset. We evaluate ReLoRA without a full-rank training warmup for a fair comparison.
For GaLore, we set subspace frequency $T$ to 200 and scale factor $\alpha$ to 0.25 across all model sizes in Table 2. For each model size, we pick the same rank $r$ for all low-rank methods, and we apply them to all multi-head attention layers and feed-forward layers in the models. We train all models using Adam optimizer with the default hyperparameters (e.g., $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$). We also estimate the memory usage based on BF16 format, including the memory for weight parameters and optimizer states. As shown in Table 2, GaLore outperforms other low-rank methods and achieves comparable performance to full-rank training. We note that for 1B model size, GaLore even outperforms full-rank baseline when $r=1024$ instead of $r=512$. Compared to LoRA and ReLoRA, GaLore requires less memory for storing model parameters and optimizer states. A detailed training setting of each model and memory estimation for each method are in the appendix.
@@ 5.2 GaLore with Memory-Efficient Optimizers
We demonstrate that GaLore can be applied to various learning algorithms, especially memory-efficient optimizers, to further reduce the memory footprint. We apply GaLore to AdamW, 8-bit Adam, and Adafactor optimizers (Shazeer & Stern, 2018; Loshchilov & Hutter, 2019; Dettmers et al., 2022). We consider Adafactor with first-order statistics to avoid performance degradation.
@@ 5.3 Scaling up to LLaMA 7B Architecture
Scaling ability to 7B models is a key factor for demonstrating if GaLore is effective for practical LLM pre-training scenarios. We evaluate GaLore on an LLaMA 7B architecture with an embedding size of 4096 and total layers of 32. We train the model for 150K steps with 19.7B tokens, using 8-node training in parallel with a total of 64 A100 GPUs. Due to computational constraints, we compare 8-bit GaLore ($r=1024$) with 8-bit Adam with a single trial without tuning the hyperparameters. As shown in Table 3, after 150K steps, 8-bit GaLore achieves a perplexity of 14.65, comparable to 8-bit Adam with a perplexity of 14.61.
@@ 5.5 Measurement of Memory and Throughput
Training 7B models on consumer GPUs with 24G memory. As shown in Fig. 4, 8-bit GaLore requires significantly less memory than BF16 baseline and 8-bit Adam, and only requires 22.0G memory to pre-train LLaMA 7B with a small per-GPU token batch size (up to 500 tokens). This memory footprint is within 24GB VRAM capacity of a single GPU such as NVIDIA RTX 4090. In addition, when activation checkpointing is enabled, per-GPU token batch size can be increased up to 4096. While the batch size is small per GPU, it can be scaled up with data parallelism, which requires much lower bandwidth for inter-GPU communication, compared to model parallelism. Therefore, it is possible that GaLore can be used for elastic training (Lin et al., 2019) 7B models on consumer GPUs such as RTX 4090s.
@@ 6 Ablation Study
How many subspaces are needed during pre-training?
We observe that both too frequent and too slow changes of subspaces hurt the convergence, as shown in Fig. 5 (left). The reason has been discussed in Sec. 4.1. In general, for small $r$, the subspace switching should happen more to avoid wasting optimization steps in the wrong subspace, while for large $r$ the gradient updates cover more subspaces, providing more cushion.
@@
How does the rank of subspace affect the convergence?
Within a certain range of rank values, decreasing the rank only slightly affects the convergence rate, causing a slowdown with a nearly linear trend. As shown in Fig. 5 (right), training with a rank of 128 using 80K steps achieves a lower loss than training with a rank of 512 using 20K steps. This shows that GaLore can be used to trade-off between memory and computational cost. In a memory-constrained scenario, reducing the rank allows us to stay within the memory budget while training for more steps to preserve the performance.
@@ 7 Conclusion
We identify several open problems for GaLore, which include (1) applying GaLore on training of various models such as vision transformers (Dosovitskiy et al., 2021) and diffusion models (Ho et al., 2020), (2) further enhancing memory efficiency by employing low-memory projection matrices, and (3) exploring the feasibility of elastic data distributed training on low-bandwidth consumer-grade hardware.
@@
We hope that our work will inspire future research on memory-efficient training from the perspective of gradient low-rank projection. We believe that GaLore will be a valuable tool for the community, enabling the training of large-scale models on consumer-grade hardware with limited resources.
Impact Statement
@@
This paper aims to improve the memory efficiency of training LLMs in order to reduce the environmental impact of LLM pre-training and fine-tuning. By enabling the training of larger models on hardware with lower memory, our approach helps to minimize energy consumption and carbon footprint associated with training LLMs.
Acknowledgments
We thank Meta AI for computational support. We appreciate the helpful feedback and discussion from Florian Schäfer, Jeremy Bernstein, and Vladislav Lialin. B. Chen greatly appreciates the support by Moffett AI. Z. Wang is in part supported by NSF Awards 2145346 (CAREER), 02133861 (DMS), 2113904 (CCSS), and the NSF AI Institute for Foundations of Machine Learning (IFML). A. Anandkumar is supported by the Bren Foundation and the Schmidt Sciences through AI 2050 senior fellow program.
Appendix A Additional Related Works
Adafactor (Shazeer & Stern, 2018) achieves sub-linear memory cost by factorizing the second-order statistics by a row-column outer product. GaLore shares similarities with Adafactor in terms of utilizing low-rank factorization to reduce memory cost, but GaLore focuses on the low-rank structure of the gradients, while Adafactor focuses on the low-rank structure of the second-order statistics.
GaLore can reduce the memory cost for both first-order and second-order statistics, and can be combined with Adafactor to achieve further memory reduction. In contrast to the previous memory-efficient optimization methods, GaLore operates independently as the optimizers directly receive the low-rank gradients without knowing their full-rank counterparts.
The fused backward operation proposed by LOMO (Lv et al., 2023b) mitigates the memory cost of storing weight gradients during training. Integrated with the standard SGD optimizer, LOMO achieves zero optimizer and gradient memory cost during training. AdaLOMO (Lv et al., 2023a) enhances this approach by combining the fused backward operation with adaptive learning rate for each parameter, similarly achieving minimal optimizer memory cost.
While LOMO and AdaLOMO represent significant advancements in memory-efficient optimization for fine-tuning or continual pre-training, they might not be directly applicable to pre-training from scratch at larger scales. For example, the vanilla Adafactor, adopted by AdaLOMO, has been demonstrated to lead to increased training instabilities at larger scales (Rae et al., 2021; Chowdhery et al., 2023; Wortsman et al., 2023; Zhai et al., 2022). We believe integrating GaLore with the fused backward operation may offer a promising avenue for achieving memory-efficient large-scale pre-training from scratch.
Appendix B Proofs
B.1 Reversibility
Definition B.1 (Reversiblity (Tian et al., 2020)).
@@ B.1 Reversibility
A network $\mathcal{N}$ that maps input ${\bm{x}}$ to output ${\bm{y}}=\mathcal{N}({\bm{x}})$ is reversible, if there exists $L({\bm{x}};W)$ so that ${\bm{y}}=L({\bm{x}};W){\bm{x}}$, and the backpropagated gradient ${\bm{g}}_{\bm{x}}$ satisfies ${\bm{g}}_{\bm{x}}=L^{\top}({\bm{x}};W){\bm{g}}_{\bm{y}}$, where ${\bm{g}}_{\bm{y}}$ is the backpropagated gradient at the output ${\bm{y}}$. Here $L({\bm{x}};W)$ depends on the input ${\bm{x}}$ and weight $W$ in the network $\mathcal{N}$.
@@
Softmax Case. Note that for softmax objective with small logits, we can also prove a similar structure of backpropagated gradient, and thus Theorem 3.2 can also apply.
Lemma B.2 (Gradient structure of softmax loss).